Solve for x,y by elimination
3x+2y=-1
6x+4y=0

Answer by stanbon(75887)

Solve for x,y by elimination
3x+2y=-1
6x+4y=0
----
Modify for elimination::
6x + 4y = -2
6x + 4y = 0
-------
Subtract to get:
0 = 2
----
That is a contradiction.
The lines are parallel.
There is no point of intersection.
There is no solution for the system of equations.
----
Cheers,
Stan H.
--------